# A transformational approach to collective behavior

### v2

@@ V Generative AI Simulation and control
Given this theory, we move on to the practical applications of simulation and control of collective systems. This application will use the concepts of Artificial Intelligence (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016), as interpreted by Glinsky (2024a, b), to approximate both the generating functional, $S_{p}[f(x)]$, and the generating function, $S_{P}(x)$. The computational pipeline to characterize and simulate the collective system is shown in Fig. 6. This pipeline contains two core components and their inverses – one to approximate the generating functional (that is, HST+PCA and iPCA+iHST), and one to approximate the generating function (that is, HJB and iHJB). There are also some practical components in the pipeline to approximate coordinate transformation functions.
@@
The phase space for this system is shown in Fig. 11. This system has two stable equilibriums, shown as the o-points, and two unstable equilibriums, shown as the x-points. One set is for the electron at 00 and $-0.3$, and another set is for the ion at $-0.9$ and $\infty$. There are basins for the GCA and the DP bound motions separated from the free motion by the black boundary called the separatrix. The o-points are stable local minimums in the energy, and the x-points are unstable saddle points that are local maximums in the energy in the vertical direction. It will take an infinite amount of time to approach the saddle points so that they will be metastable. An external thermal force coming from a heat bath will exert a force perpendicular to the motion towards lower energy. This will cause the motion to descend from the mountain top, spiralling around the mountain until the mountain pass, that is the saddle point, is reached. It will take a long time to reach the mountain pass, and an equally long time to move away from the mountain pass. The motion will then fall into one of the two basins of bound motion and spiral down to the valley center, that is the stable equilibrium. The motion will eventually wander around the valley center in thermal equilibrium
@@
where $\omega_{0}\ll\omega_{\text{sp}}$, $J_{0}\,\omega_{0}\lesssim f_{0}\ll J_{0}\,\omega_{\text{sp}}$ (so that the ponderomotive force is large but the motion is small), $J_{0}\equiv E_{0}/\omega_{0}$, and $\varepsilon_{P}$ is a random $\Delta P$ of size $\epsilon_{P}\ll J_{0}$ taken every $2\pi/\omega_{0}$. This $F_{\text{sp}}(\tau)$ force is not dependant on $P^{*}$ or $P$, just the time invariant mapping generated by $S_{P}(q)$, that is $p(P,Q)$ and $q(P,Q)$, and the functional transformation iPCA+iHST generated by $S_{p}[f(x)]$, that is $f[p(\tau),q(\tau)](x)$ and $\pi[p(\tau),q(\tau)](x)$. This pipeline is vibrating the collective system with a high frequency carrier signal modulated with the characteristic “local spectrums” of the collective system, that is $\beta^{*}_{i}(z)$. This is a set of measure zero compared to the Hilbert space of possible “local spectrums”.
@@
More details on the mathematics and concepts presented in this section, as well as other ways of using them to control the system, can be found in Glinsky and Sievert (2023) and Glinsky (2024c).
@@ VIII Discussion and Conclusions
The concepts of genAI (Glinsky, 2024a) are utilized to approximate the generator of the individual motion, $S_{P}(q)$, and the generator of the collective motion, $S_{p}[f(x)]$. In Sec. V, a computational pipeline was designed using architectures for the $S_{p}[f(x)]$ (that is the HST, shown in Fig. 4) and for the $S_{P}(q)$ (that is the HJB, shown in Fig. 7) that can be trained with much less effort than the traditional approaches of DRL and GPTs.
The traditional approaches to genAI are closely related to the computational pipeline that we designed. Optimal control, as done in genAI with Deep Q-Learning (DQN) (Mnih et al., 2015), also known as DRL, is based on parametric estimations of a Q-function
@@
where $s=q$ is the state, $a=p$ are the actions to be taken or the co-state, $\theta=P$ are the parameters of the estimator, and the approximate Q-function is the action. Meanwhile, the generator of the flow, as done in genAI with Generative Pretrained Transformers (GPTs), is based on parametric estimation of a log-likelihood
@@
where $x=q$ is the coordinate, $\theta=P$ are the parameters of the estimator, and the approximate log-likelihood is the action.

### v6

@@ V Generative AI Simulation and control
Given this theory, we move on to the practical applications of simulation and control of collective systems. This application will use the concepts of Artificial Intelligence (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016), as interpreted by Glinsky (2024b, c), to approximate both the generating functional, $S_{p}[f(x)]$, and the generating function, $S_{P}(x)$. The computational pipeline to characterize and simulate the collective system is shown in Fig. 7. This pipeline contains two core components and their inverses – one to approximate the generating functional (that is, HST+PCA and iPCA+iHST), and one to approximate the generating function (that is, HJB and iHJB). There are also some practical components in the pipeline to approximate coordinate transformation functions.
@@
The phase space for this system is shown in Fig. 12. This system has two stable equilibriums, shown as the o-points, and two unstable equilibriums, shown as the x-points. One set is for the electron at 00 and $-0.3$, and another set is for the ion at $-0.9$ and $\infty$. There are basins for the GCA and the DP bound motions separated from the free motion by the black boundary called the separatrix. The o-points are stable local minimums in the energy, and the x-points are unstable saddle points that are local maximums in the energy in the vertical direction. It will take an infinite amount of time to approach the saddle points so that they will be metastable. An external thermal force coming from a heat bath will exert a force perpendicular to the motion towards lower energy. This will cause the motion to descend from the mountain top, spiraling around the mountain until the mountain pass, that is the saddle point, is reached. It will take a long time to reach the mountain pass, and an equally long time to move away from the mountain pass. The motion will then fall into one of the two basins of bound motion and spiral down to the valley center, that is the stable equilibrium. The motion will eventually wander around the valley center in thermal equilibrium
@@
where $\omega_{0}\ll\omega_{\text{sp}}$, $J_{0}\,\omega_{0}\lesssim f_{0}\ll J_{0}\,\omega_{\text{sp}}$ (so that the ponderomotive force is large but the motion is small), $J_{0}\equiv E_{0}/\omega_{0}$, and $\varepsilon_{P}$ is a random $\Delta P$ of size $\epsilon_{P}\ll J_{0}$ taken every $2\pi/\omega_{0}$. This $F_{\text{sp}}(\tau)$ force is not dependent on $P^{*}$ or $P$, just the time invariant mapping generated by $S_{P}(q)$, that is $p(P,Q)$ and $q(P,Q)$, and the functional transformation iPCA+iHST generated by $S_{p}[f(x)]$, that is $f[p(\tau),q(\tau)](x)$ and $\pi[p(\tau),q(\tau)](x)$. This pipeline is vibrating the collective system with a high frequency carrier signal modulated with the characteristic “local spectrums” of the collective system, that is $\beta^{*}_{i}(z)$. This is a set of measure zero compared to the Hilbert space of possible “local spectrums”.
@@
More details on the mathematics and concepts presented in this section, as well as other ways of using them to control the system, can be found in Glinsky and Sievert (2023) and Glinsky (2024d).
@@ VIII Discussion and Conclusions
The concepts of genAI (Glinsky, 2024b) are utilized to approximate the generator of the individual motion, $S_{P}(q)$, and the generator of the collective motion, $S_{p}[f(x)]$. In Sec. V, a computational pipeline was designed using architectures for the $S_{p}[f(x)]$ (that is the HST, shown in Fig. 4) and for the $S_{P}(q)$ (that is the HJB, shown in Fig. 8) that can be trained with much less effort than the traditional approaches of DRL and GPTs.
The traditional approaches to genAI are closely related to the computational pipeline that we designed. Optimal control, as done in genAI with Deep Q-Learning (DQN) (Mnih et al., 2015), also known as DRL, is based on parametric estimations of a Q-function or value function
@@
where $s=q$ is the state, $a=p$ are the actions to be taken or the co-state, $\theta=P$ are the parameters of the estimator, and the approximate Q-function is the action. Meanwhile, the generator of the flow, as done in genAI with Generative Pretrained Transformers (GPTs), is based on parametric estimation of a log-likelihood or score function
@@
where $x=q$ is the coordinate, $\theta=P$ are the parameters of the estimator, and the approximate log-likelihood is the action. When it comes to estimating the probability distributions, that is treating the system statistically or quantizing the theory, our work considers a dynamic equilibrium that is uniform in $Q$, and is an arbitrary function of $P$, such that the probability is
$\rho(P)\,\text{d}P\,\text{d}Q,$
where $\rho(P)$ is an arbitrary function. In contrast, the methods of genAI consider a statistical equilibrium that is not only uniform in $Q$, but also a thermal distribution in $P$, such that the probability is
$\text{e}^{-E(P)/k_{\text{B}}T}\,\text{d}P\,\text{d}Q.$
This is true of all of the current methods of genAI, whether that be DQNs, GPTs, or Boltzmann Machines as estimated by Hopfield Networks, also known as MLPs (Hopfield, 1982; Ackley et al., 1985).
@@
With respect to the relevance and applicability of conservative systems, two things should be noted. Application of an external force, like a dissipative frictional force, or a diffusive Fokker-Planck thermal force or a general Boltzmann collisional force, to a conservative system does not make the system non-conservative or even change the system. It only changes the energy of the system, that is $\Delta E=E(P^{\prime})-E(P)$, where $P^{\prime}=P+\Delta P$ and $\Delta P$ is the change in action. Application of a conservative force, like a potential barrier, to a conservative system does not make the system non-conservative. It also only changes the energy of the system, that is $\Delta E=E^{\prime}(P)-E(P)$, where $E^{\prime}=E+V$ and $V$ is the added potential energy.